OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = x2 - 2x + 1, what is f(g(x)) ?

OpenStudy (anonymous):

4*(x2 - 2x + 1) + 1

OpenStudy (shayaan_mustafa):

\[\large f(g(x))=4(x ^{2}-2x+1)+1\]\[\large f(g(x))=4x ^{2}-8x+4+1\]\[\large f(g(x))=4x ^{2}-8x+5\]